OpenStudy (anonymous):

2. The ratio of clown fish to angel fish at a pet store is 5:4. The ratio of angel fish to goldfish is 4:3. There are 60 clown fish at the pet store. a. How many angel fish are there? b. How many goldfish are there?

jimthompson5910 (jim_thompson5910):

"The ratio of clown fish to angel fish at a pet store is 5:4" "There are 60 clown fish at the pet store." those two facts allow us to make this equation 5/4 = 60/x where x is the number of angel fish

jimthompson5910 (jim_thompson5910):

5/4 = 60/x 5*x = 4*60 5x = 240 x = ???

OpenStudy (anonymous):

x=48

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

do you see how to do part B?

OpenStudy (anonymous):

yes thank you jim

jimthompson5910 (jim_thompson5910):

yw